OpenStudy (anonymous):

Solve for x: 1 over 4 (2x − 14) = 4

OpenStudy (anonymous):

\[1/4*(2x-14)=4\] Is this the equation you want me to solve for? If so: 1/4*(2x-14)=4 2x-14=16 2x=30 x=15